Solution A is 25% acid, and solution B is 40% acid. How much of each is needed to make 60L of a solution that is 30% acid?
1
Expert's answer
2020-04-24T18:29:03-0400

Since we are not given the density of solutions, let density of all solutions = 1 g/L. Let the mass of the solution A = x kg, then the mass of the solution B = 60-x kg. Knowing the mass fraction of acid in all solutions, we make the equation with one unknown:


0.25x + (60-x)0.4 = 60*0.3

0.25x + 24 - 0.4x = 18

0.15x = 6

x = 40 kg


So, mass of the solution A = x = 40 kg, mass of the solution B = 60 - 40 = 20 kg.